OpenStudy (anonymous):

(1+cosx)(1+secx)=sinxtanx prove the following trig identity.

mathslover (mathslover):

Hello @rl88

OpenStudy (anonymous):

Hey!

OpenStudy (anonymous):

Help pleasee! :D

mathslover (mathslover):

Did you try to do that first?

OpenStudy (anonymous):

\[(1+a)(1+\frac{1}{a})\] is the first step

mathslover (mathslover):

or I will say that distributive property is the first hint.

mathslover (mathslover):

@rl88 do you know distributive property?

OpenStudy (anonymous):

what is distributive property?

mathslover (mathslover):

(a+b)(c+d) = ?

OpenStudy (anonymous):

actually, now that i look more carefully, i am not sure it is true

mathslover (mathslover):

@satellite73 will help you @rl88 I am not feeling well. Also welcome to openstudy @rl88

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

oh maybe it is true, hmm

OpenStudy (anonymous):

i'm confused...

OpenStudy (anonymous):

i think this is false is what i am saying

OpenStudy (anonymous):

suppose \(x=0\) then the left hand side is \(1+\cos(0))(1+\sec(0))=(1+1)(1+1)=4\) but the right hand side is \(\sin(0)\tan(0)=0\times 0=0\) so this is not an identity

OpenStudy (anonymous):

look carefully at the problem and make sure you copied it correctly, because this one is not true

OpenStudy (anonymous):

hmm no, its right, theres another question for example 1-2sin^2x=2cos^2x-1 the rearrangement of that question gets you a true statement but that statement alone doesn't if you substitute x for 0.

OpenStudy (anonymous):

an "identity" means true for all values of \(x\) like \(2(x-4)=2x-8\) is an identity it cannot only be true sometimes

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

if you graph both of them above, you will see that they are not the same, but i any case you do not need to do that, because letting \(x=0\) and getting \(4=0\) proves that it is not an identity

OpenStudy (anonymous):

hmm...OK thanks!

OpenStudy (anonymous):

i know what it is supposed to be i think

OpenStudy (anonymous):

\((1-\cos(x))(1+\sec(x))\) might be right

OpenStudy (anonymous):

maybe that was the actual problem, show \[(1-\cos(x))(1+\sec(x))=\sin(x)\tan(x)\] that i think we can do

OpenStudy (anonymous):

might that have been the question?

OpenStudy (anonymous):

maybe it was a typo on my page? it could have been? how would that be done?

OpenStudy (anonymous):

it is easier to write the algebra if you replace \(\cos(x)\) by \(a\) you get \[(1-a)(1+\frac{1}{a})=1+\frac{1}{a}-a-1=\frac{1}{a}-a=\frac{1-a^2}{a}\]

OpenStudy (anonymous):

then since \(a=\cos(x)\) the numerator is \(1-\cos^2(x)=\sin^2(x)\) making the whole fraction \[\frac{\sin^2(x)}{\cos(x)}=\sin(x)\tan(x)\]

OpenStudy (anonymous):

Hey is anyone here good in English and knows a lot about the book Hamlet please hmu with a message! =) ty

OpenStudy (anonymous):

ohh...i dont get the last bit? how does sin^2x/cosx =sinxtanx?